OpenStudy (anonymous):

Can anyone help me with any of these math problems 1) Write the algebraic expression corresponding to: The quotient of x decreased by 5 and the quantity twice y increased by 7. perform the operation and simplify (for problems involving mixed numbers convert your answer to a mixed number)

OpenStudy (anonymous):

2)\[\frac{ 3 }{ 5 }.\frac{ 5 }{ 7} =\]

OpenStudy (anonymous):

\[3) \frac{ 5 }{ 21 } \div \frac{ 25 }{ 14}\]